OpenStudy (anonymous):

Jolene invests her savings in two bank accounts, one paying 4 percent and the other paying 9 percent simple interest per year. She puts twice as much in the lower-yielding account because it is less risky. Her annual interest is 3230 dollars. How much did she invest at each rate? Amount invested at 4 percent interest is $ ?? Amount invested at 9 percent interest is $ ?? I need help

OpenStudy (stamp):

0.04(2a)+0.09(a)=3230 0.08a + 0.09a = 3230 0.17a = 3230 a = 3230*100/17

OpenStudy (stamp):

a = $19,000 2a = $38,000

OpenStudy (stamp):

\[4\%($38,000)+9\%($19,000)=\]\[$1520+$1710=\Sigma_i\]\[$3230=\Sigma_i\]It checks out.

OpenStudy (anonymous):

so what is invested at 4 percent

OpenStudy (stamp):

@gil714 I just did the whole problem for you. Read the problem, then read the problem again, and after you finish reading the problem, read the problem. Then you will likely have found the answer to your question.